find the maximum and minimum value of g(x) = x^3 + 1
plz reply as fast as you can
 

Grade:12

2 Answers

Vikas TU
7 years ago
 g(x) = x^3 + 1
differentiate it 
g’(x) = 3x^2 = 0
hence slope is always positive.
the graph will approach to + infinity  and   – infinity
as x increases and decreases.
 
Therfoore maximum and minmum values of g(x) is:
+infinity  and  -infinity  respectively.
as x has no boundation or we can say that
x belongs to R.
 
SREEKANTH
7 years ago
the given function is g(x)=x^3+1
for determining maximum and minimum values we have to find out that the given function is increasing or decreasing so differentiate the given function on differentiating we get
 d/dx(x^3+1)=3x^2  which is alway increasing and continous function for all x so the maximum and minimum values are +infinity and -infinity